# Establishing a Connection Between the Jet and the Corona in Black Hole Low Mass X-Ray Binaries

### v1

@@ 3 Results
Our X-ray spectral modeling aimed to constrain the height of the corona above the accretion disk and examine its correlation with the jet’s behavior. Figures 1, 2, and 3 show the variation of the coronal height and with time. In all three sources, the corona’s height maintains a low, nearly constant value during the HS (see Table LABEL:results_table). However, around the time when the compact jet is quenched, and a transient jet is launched, the corona height increases markedly.
@@
The results for J1820 are visualized in Figure 1. During its long initial HS, the corona height hovers at relatively low values, between and until MJD 58259 when it jumps up to before returning to by MJD 58297. Following this, the inferred height of the corona increases to on MJD 58306. This happens concurrently with the radio spectrum becoming more optically thin as decreases to on MJD 58305, then to on MJD 58311. These changes also coincide with the jet ejection event on MJD 58306 (Wood et al., 2021). The corona height remains high for the following two epochs before moving back toward the black hole as the SS progresses. During the second IS, the compact jet is detected again, and corona height reaches . In the final HS, as recovers, the corona height returns to low values of about .
@@
A similar pattern is observed in J1348 (Figure 2). Our observations begin in the initial HS with the corona height low ( ). The corona height subsequently increases to . on MJD 58521.07 The increase in coronal height again coincides with ’s decrease from low positive to steep negative values and the ejection of the jet (Carotenuto et al., 2021). The following observations see the corona return to lower values. In this source, the corona height is modeled at during MJD 58577.03; this immediately precedes the projected date of a second ejection of plasma blob (Carotenuto et al., 2021). As returns to positive values during the final HS, we measure low coronal heights again.
@@
J1535 (Figure 3) also exhibits a low corona height during the initial HS. The height quickly increases to a maximum height of as decreases to negative values (Russell et al., 2019). This increase in coronal height again occurs around the time of the jet ejection (Chauhan et al., 2021; Russell et al., 2019). The corona height falls to low values again for the remainder of our epochs.
@@ 4 Discussion
Our results suggest a causal relationship between the jet and the corona, revealing a change in the corona geometry contemporaneous with a jet ejection event. As outlined in Section 3, the main outcomes of our work are the following:
@@
1) In the initial HSs, we observed that the X-ray corona remains steadily close to the black hole, and the compact radio jet is also observed to remain mostly unchanged.
@@
2) As the LMXB transitions from the HS to the SS, we found that the corona height increases dramatically. At around the same time, the compact jet emission is quenched and jet ejecta are launched from the system. This jet change is determined from spatially resolved ejecta and a transition to optically-thin synchrotron emission, as emitted by the transient jet.
@@
3) In J1820, as the outburst fades and the source returns to a hard state, the compact jet recovers (Echiburú-Trujillo et al., 2024), the corona height once again increases before returning to the its original low levels at later times.
@@
The idea of a connection between the radio jet and the corona is not new; for over two decades it has been postulated that the jet or jet base is connected to the corona and responsible for the relativistic reflection features in X-ray binaries (Markoff et al., 2001b, 2003; Markoff & Nowak, 2004). In a model proposed by Markoff et al. (2001b), it has been argued that inverse Compton and synchrotron emission from the base of the jet could be appreciably contributing to the non-thermal power-law spectra or even produce the detected X-ray reflection (Markoff & Nowak, 2004).
@@
Various X-ray spectral and timing studies have already noted that the corona is dynamic in nature, i.e., the corona geometry changes with the accretion rate change (e.g., Kara et al., 2019; García et al., 2020; Lucchini et al., 2021). However, the exact nature of this evolution remains unclear. Kara et al. (2019) studied reverberation lags in J1820 during its initial HS and found that the corona gradually contracted as the source softened. Lucchini et al. (2021) further strengthened this notion and suggested a link between jet-base and corona. In MAXI J1836–194, they found that harder states () require a more compact corona/jet-base region close to the black hole, while softer states require acceleration to move downstream. Further studies of quasi-periodic oscillations evolution led some to suggest a dual corona model (García et al., 2020), that is sometimes also seen as a vertically extended corona/jet ecosystem that varies in size (Zhang et al., 2023; Sripada & Kandulapati, 2022; Liao et al., 2024; Liu et al., 2022; García et al., 2022). To account for the effects of an extended corona on the observed spectrum, Lucchini et al. (2023) invoked a dual corona model in which a lower component (several ) dominates the features observed in the time-average spectrum, while the second component, a higher corona (hundreds of ) primarily accounts for the time-variable features. Lucchini et al. (2023) tested this model on J1820 (NICER observation on MJD 58301). The authors found a lower corona at height , and a much fainter upper corona at height . That NICER observation occurs between our observations on MJDs 58297 and 58306, where our single-lamppost corona height went from to . Our simpler model is consistent with these results if our single lamp-post corona height can be seen as a weighted average of their two heights. The success of their model in explaining the both the time-variability and the time-averaged X-ray spectrum points to a likely scenario: that while a core of the corona resides somewhat close to the black hole, elements of it stretch out to greater distances. Both their model and ours act as probes of this more extended corona.
@@
During the HS, the corona height always remained at a low value or nearly constant (see Table LABEL:results_table), and the radio spectrum is consistently flat or inverted. This suggests that the geometry of the corona is close to the black hole at this time. Russell et al. (2020) found that the compact particle-acceleration region of J1535 lies at distances of thousands of from the black hole during the HS. Echiburú-Trujillo et al. (2024) found that this region lies even further away during J1820’s HS. This indicates that the corona and the jet particle acceleration region are two separate entities.
@@
However, we observed an increase in the corona height around the time of the transient jet ejection, suggesting a connection between the two. We note that the transient jet is launched around the same time as the compact jet quenching in our three sources; as such, due to the low cadence of our X-ray data, we cannot precisely identify what the immediate connection is. We do see an increase in the coronal height at the same time as the compact jet is re-established in J1820, suggesting that the connection may be with the compact jet. Either way, this work implies that the corona and the outflow are connected through some physical process. Considering our findings and following the models of Markoff et al. (2001b); Markoff & Nowak (2004), we argue that the corona and the jet acceleration region are coupled, with continuous feedback between them (Malzac, 2014). A magnetic field likely originating from the accretion disk, or threaded in the event horizon could be the mechanism behind this coupling. Parts of the corona could be ejected out far beyond the lamp-post height assumptions and possibly interact with the first acceleration region or jet base through the magnetic field lines (Narayan et al., 2003; Igumenshchev, 2008; Tchekhovskoy et al., 2011; Janiuk & James, 2022).
@@
An alternative explanation could be that a change in the accretion rate close to the HS to SS transition (point of transient jet eject) makes fast-moving matter clump into slow-moving plasma, creating a shock in the accretion disk (Fender et al., 2004c; Narayan et al., 2012; Janiuk & James, 2022). This shock changes the accompanying magnetic field configuration (Narayan et al., 2003; Igumenshchev, 2008; Tchekhovskoy et al., 2011; Janiuk & James, 2022), and leads to a change in the corona geometry, possibly from compact to the lamp-post structure, which leads to a change in the coronal height.
@@
This also suggests that the corona might have a more complex geometry, and it could be a combination of both an extended corona and the near-point source. Rapid changes to the coronal geometry as the jet transitions to transient have already been observed in Swift J1727.8–1613 (Podgorný et al., 2024). The authors also reported a return to pre-outburst polarization levels upon re-entering the HS, indicating that the coronal geometry and nature were similar to those before the outburst (Podgorný et al., 2024). Furthermore, a drastic decrease in polarization during the SS has been widely observed (Mastroserio et al., 2024; Svoboda et al., 2024). The polarization findings also lend support to the notion of a non-trivial coronal geometry; a radially extended portion is especially indicated by polarization (Veledina et al., 2023; Krawczynski et al., 2022; Ingram et al., 2024; Steiner et al., 2024).
@@
Our analysis cannot determine whether the jet ejection or the coronal height increase occurred first or whether the corona is entwined with the compact jet or the transient jet. Jet ejection times are difficult to pinpoint (Wood et al., 2021; Russell et al., 2020; Carotenuto et al., 2021). Additionally, our X-ray spectral epochs are too irregular to track the exact onset of coronal height changes, making it challenging to identify the cause of these geometric changes. Higher-cadence simultaneous observations in radio and X-ray bands could clarify whether the coronal height increase precedes jet ejection, providing new insights.
@@ Acknowledgements
This work is partly supported by the NSF REU Program (Award Number 2244344) at Montana State University. This work is additionally supported by the Montana Space Grant Consortium (Grant Number 80NSSC20M0042). JC and RS acknowledge support from the Leverhulme Trust grant RPG-2023-240.

### v2

@@ 3 Results
Our X-ray spectral modeling aimed to constrain the height of the corona above the accretion disk and examine its correlation with the jet’s behavior. Figures 1, 2, and 3 show the variation of the coronal height and with time. In all three sources, the corona’s height maintains a low, nearly constant value during the LHS (see Table 1). However, around the time when the compact jet is quenched, and a transient jet is launched, the corona height increases markedly.
@@
The results for J1820 are visualized in Figure 1. During its long initial LHS, the corona height hovers at relatively low values, between and until MJD 58259 when it jumps up to before returning to by MJD 58297. Following this, the inferred height of the corona increases to on MJD 58306. This happens concurrently with the radio spectrum becoming more optically thin as decreases to on MJD 58305, then to on MJD 58311. These changes also coincide with the jet ejection event on MJD 58306 (Wood_2021). The corona height remains high for the following two epochs before moving back toward the black hole as the HSS progresses. During the transition to the second LHS, the compact jet is detected again, and corona height reaches . In the final LHS, as recovers, the corona height returns to low values of about .
@@
A similar pattern is observed in J1348 (Figure 2). Our observations begin in the initial LHS with the corona height low ( ). The corona height subsequently increases to . on MJD 58521.07 The increase in coronal height again coincides with ’s decrease from low positive to steep negative values and the ejection of the jet (car). The following observations see the corona return to lower values. In this source, the corona height is modeled at during MJD 58577.03; this immediately precedes the projected date of a second ejection of plasma blob (car). As returns to positive values during the final LHS, we measure low coronal heights again.
@@
J1535 (Figure 3) also exhibits a low corona height during the initial LHS. The height quickly increases to a maximum height of as decreases to negative values (Russell_2019). This increase in coronal height again occurs around the time of the jet ejection (chauhan_2021; Russell_2019). The corona height falls to low values again for the remainder of our epochs.
In J1820 and J1535, the decision to separate our epochs into groups led to a different set of “global” parameters for each group. There are notable variations in some of these parameters between different spectral groupings of the same source. Most notable of these differences is the gradual change in inclination through the three groups of J1820 from 26∘ in group 1 to 30∘ in group 2 to 47∘ in group 3. Fitting was initially performed with the inclination and other global parameters of groups 1 and 3 frozen to best fit values of the those in group 2. This resulted in a poorer , but still preserved our reported trend in coronal height changes. While significant changes to the inclination of LMXB GRS 1915+105 have been observed to occur within a year (Rodriguez_2025), this is not something generally observed or expected in LMXBs. Another global parameter exhibiting changes between groups is the tbabs equivalent hydrogen column density parameter nH, which goes from about 0.4 atoms cm-2 in group 1 to about 0.07 atoms cm-2 in groups 2 and 3 of J1820. A version of our J1820 group 1 fits with the nH parameter fixed to 0.07 was performed with results shown in the appendix alongside our reported results. This resulted in a poorer , with no significant change to the corona heights. We note that the number of epochs using supplementary NICER or Swift/XRT data differs between groups, potentially contributing to differences in some of the global parameters. However, this is unlikely to affect the observed trend in corona heights, as illustrated by our attempts to mitigate the global parameter differences. Additionally, the major increases and decreases in corona height are seen to occur within our individual spectral groups.
@@ 4 Discussion
Our results suggest a relationship between the jet and the corona, revealing a change in the corona geometry contemporaneous with a jet ejection event. As outlined in section 3, the main outcomes of our work are the following:
@@
1) In the initial LHSs, we observed that the X-ray corona remains steadily close to the black hole, and the compact radio jet is also observed to remain mostly unchanged.
@@
2) As the LMXB transitions from the LHS to the HSS, we find that the corona height increases dramatically. Concurrently, the compact jet emission is quenched, and transient jet ejecta are launched from the system. This change in jet behavior is identified through spatially resolved ejecta and the emergence of optically thin synchrotron emission, characteristic of the short-lived transient jet before the jet disappears entirely as the source enters the HSS.
@@
3) In J1820, as the outburst fades and the source returns to a hard state, the compact jet recovers (Echibur_Trujillo_2024), the corona height once again increases before returning to the its original low levels at later times.
@@
The idea of a connection between the radio jet and the corona is not new; for over two decades it has been postulated that the jet or jet base is connected to the corona and responsible for the relativistic reflection features in X-ray binaries (Markoff2001; Markoff2003; Markoff2004). In a model proposed by Markoff2001, it has been argued that inverse Compton and synchrotron emission from the base of the jet could be contributing to the non-thermal power-law spectra or even produce the detected X-ray reflection (Markoff2004). Alternatively, it has been proposed that the Compton-scattering of soft photons could occur entirely within the jet during the LHS (Reig2003; Giannios2004; kylafis2008), and that the lingering power-law component seen during the HSS could be produced by a cool disk extending all the way to the ISCO rather than a corona (Reig2001). Comptonization by the jet rather than a corona has been somewhat successful in explaining the reflection component (Reig2021), however this model predicts a flatter emissivity than what reflection model fits typically find (uttley2025). Wang_2020hh suggested that there are two Comptonization regions, with one being the more compact corona near the black hole suggested by kara19, and the other occurring within the jet.
Various X-ray spectral and timing studies have already noted that the corona is dynamic in nature, i.e., the corona geometry changes with the accretion rate change (e.g., kara19; Garc_a_2020; Lucchini2021; alabarta2025; rawat2023; ma2023). Studies of QPOs in LMXBs have revealed much about the dynamics of the corona. Fast variations in the QPO time lags of GRS 1915+105 suggest that the size of the corona can vary by nearly an order of magnitude over the course of a single 10.5 hour observation (belloni2024). However, the exact nature of the corona evolution remains unclear. kara19 studied reverberation lags in J1820 during its initial LHS and found that the corona gradually contracted as the spectrum softened. We find that our results do not align with those reported by kara19. This discrepancy may arise because the two studies are probing different regions of the corona. While kara19 observed significant variability of the lags, which they interpreted as reverberation lags, the iron K emission line profile—on which our modeling is based—remains relatively stable within NICER’s energy band.
Lucchini2021 provided further evidence for a dynamic corona and suggested a link to the jet-base. In MAXI J1836–194, they found that harder states () require a more compact corona/jet-base region close to the black hole, while softer states require acceleration to move downstream. Further studies of quasi-periodic oscillations evolution led some to suggest a dual corona model (Garcia2021), that is sometimes also seen as a vertically extended corona/jet ecosystem that varies in size. (zhang2023; sripada2022; liao2024; Liu_2022; Garcia2022; rawat2023; Mendez2022; ma2023). For example, zhang2023 found a vertically extended corona varying from 100 to 700 during the HIMS-SIMS transition of J1535 (assuming a BH). liao2024 found corona sizes that varied between and in Swift J1727.8–1613 (for a BH of size (sanchez2025)). In a NICER spectral timing study of J1820 wang2021 found increasing soft reverberation lags during the hard-to-soft transition, a change they attribute to the corona expanding from to . This trend of an expanding corona is consistent with our results from the same time period, and they suggest this expansion could result in the launching of a jet knot. Alternatively, these increasing lags could also be understood in the context of an evolving accretion disk atmosphere, although the thermalization timescale is often very small (salveson2022)
@@
To account for the effects of an extended corona on the observed spectrum, Lucchini2023 utilized a dual corona model, originally proposed by Garcia2021, in which a lower component (several ) dominates the features observed in the time-average spectrum, while the second component, a higher corona (hundreds of ) primarily accounts for the time-variable features. Lucchini2023 tested this model on J1820 (NICER observation on MJD 58301). The authors found a lower corona at height , and a much fainter upper corona at height . That NICER observation occurs between our observations on MJDs 58297 and 58306, where our single-lamppost corona height went from to . Our simpler model is consistent with these results if our single lamp-post corona height can be seen as an flux-weighted average position of the two coronal positions. The success of their model in explaining both the time-variability and the time-averaged X-ray spectrum points to a likely scenario: that while a core of the corona resides somewhat close to the black hole, elements of it stretch out to greater distances. Both their model and ours act as probes of this more extended corona.
@@
During the LHS, the corona height always remained at a low value or nearly constant (see Table 1), and the radio spectrum is consistently flat or inverted. This suggests that the geometry of the corona is close to the black hole at this time. Russell_2020 found that the compact particle-acceleration region of J1535 lies at distances of thousands of from the black hole during the LHS. Echibur_Trujillo_2024 found that this region lies even further away during J1820’s LHS. This indicates that the corona and the jet particle acceleration region are two separate entities.
@@
However, we observed an increase in the corona height around the time of the transient jet ejection, suggesting a connection between the two. We note that the transient jet is launched around the same time as the compact jet quenching in our three sources; as such, due to the low cadence of our X-ray data, we cannot precisely identify what the immediate connection is. We do see an increase in the coronal height at the same time as the compact jet is re-established in J1820, suggesting that the connection may be with the compact jet. Either way, this work implies that the corona and the outflow are connected through some physical process. Considering our findings and following the models of Markoff2001; Markoff2004, we argue that the corona and the jet acceleration region are coupled, with continuous feedback between them (Malzac2014). A magnetic field likely originating from the accretion disk, or threaded in the event horizon could be the mechanism behind this coupling. Parts of the corona could be ejected out far beyond the lamp-post height assumptions and possibly interact with the first acceleration region or jet base through the magnetic field lines (Narayan2003; Igumenshchev2008; Tchekhovskoy2011; Janiuk2022).
@@
An alternative explanation could be that a change in the accretion rate close to the LHS to HSS transition (point of transient jet eject) makes fast-moving matter clump into slow-moving plasma, creating a shock in the accretion disk (Fender2004a; Narayan2012). This shock may change the accompanying magnetic field configuration and lead to a change in the corona geometry, possibly from compact to the lamp-post structure, which leads to a change in the coronal height.
@@
This also suggests that the corona might have a more complex geometry, and it could be a combination of both an extended corona and the near-point source. Rapid changes to the coronal geometry as the jet becomes transient have already been observed in Swift J1727.8–1613 (Podgorny2024). The authors also reported a return to pre-outburst polarization levels upon re-entering the LHS, indicating that the coronal geometry and nature were similar to those before the outburst (Podgorny2024). Furthermore, a drastic decrease in polarization during the HSS has been observed (Mastroserio2024; Soboda2024). The polarization findings also lend support to the notion of a non-trivial coronal geometry; a radially extended portion is especially indicated by polarization (Veledina_2023_ixpe; cyg_ixpe; Ingram_2024; Steiner2024) However, it should be noted that interpretation of the polarization measurements hinges on simulation-based models that only consider certain geometries. Although some reverberation studies study suggest a small corona, others have also suggested the possibility of a corona that is radially extended along the disk (ma2023; alabarta2025). Reconciling these findings with spectral modeling results suggesting a more compact or vertically extended coronal geometry is challenging. It has been suggested (e.g. alabarta2025) that the corona both begins and ends an outburst with a compact component and a large, horizontally extended component, while the large component takes on a jet-like shape at the peak of the outburst.
@@
Our analysis cannot determine whether the jet ejection or the coronal height increase occurred first or whether the corona is entwined with the compact jet or the transient jet. Jet ejection times are difficult to pinpoint (Wood_2021; Russell_2020; car). Additionally, our X-ray spectral epochs are too irregular to track the exact onset of coronal height changes, making it challenging to identify the cause of these geometric changes. Higher-cadence simultaneous observations in radio and X-ray bands could clarify whether the coronal height increase precedes jet ejection, providing new insights.
@@ Acknowledgements
We sincerely thank the referee for their thorough and constructive review. We appreciate the time and effort they put into evaluating our manuscript, and we have carefully considered all comments and suggestions. This work is partly supported by the NSF REU Program (Award Number 2244344) at Montana State University. This work is additionally supported by the Montana Space Grant Consortium (Grant Number 80NSSC20M0042). JC and RS acknowledge support from the Leverhulme Trust grant RPG-2023-240.
| MAXI J1348–630 | ||||||
| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data 333Ni=NICER, S = Swift/XRT | Supp. ObsID | Corona Heights ()444p means model pegged height at minimum value,-p means lower limit is minimum value |
| 1 | 58515.13 | LHS | 80402315002 | - | ||
| 2 | 58515.60 | LHS | 80402315004 | - | ||
| 3 | 58521.07 | HIMS | 80402315006 | S | 00011107004 | |
| 4 | 58525.26 | SIMS | 80402315008 | - | ||
| 5 | 58550.64 | HSS | 80402315010 | S | 00088843002 | |
| 6 | 58577.03 | HSS | 80402315012 | Ni | 2200530118 | |
| 7 | 58655.71 | LHS | 80502304002 | Ni | 2200530161 | |
| 8 | 58660.50 | LHS | 80502304004 | - | ||
| 9 | 58672.81 | LHS | 80502304006 | Ni | 2200530169 | |
| : | 1.07 | |||||
| MAXI J1820+070 Group 1 | ||||||
| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
| 1 | 58191.94 | LHS | 90401309002 | - | ||
| 2 | 58198.04 | LHS | 90401309004 | - | ||
| 3 | 58198.39 | LHS | 90401309006 | Ni | 1200120106 | |
| 4 | 58201.53 | LHS | 90401309008 | Ni | 1200120109 | |
| 5 | 58201.86 | LHS | 90401309010 | Ni | 1200120109 | |
| 6 | 58212.22 | LHS | 90401309012 | Ni | 1200120120 | |
| 7 | 58224.96 | LHS | 90401309013 | Ni | 1200120130 | |
| 8 | 58225.30 | LHS | 90401309014 | Ni | 1200120131 | |
| : | 1.01 | |||||
| MAXI J1820+070 Group 2 | ||||||
| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
| 9 | 58242.27 | LHS | 90401309016 | Ni | 1200120144 | |
| 10 | 58255.15 | LHS | 90401309018 | - | ||
| 11 | 58255.91 | LHS | 90401309019 | S | 00088657004 | |
| 12 | 58259.33 | LHS | 90401324002 | Ni | 1200120155 | |
| 13 | 58297.20 | LHS | 90401309021 | Ni | 1200120189 | |
| 14 | 58306.38 | SIMS | 90401309023 | Ni | 1200120198 | |
| 15 | 58315.02 | HSS | 90401309025 | Ni | 1200120208 | |
| 16 | 58327.04 | HSS | 90401309026 | - | ||
| 17 | 58327.05 | HSS | 90401309027 | - | ||
| : | 0.98 | |||||
| MAXI J1820+070 Group 3 | ||||||
| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
| 18 | 58343.61 | HSS | 90401309029 | - | ||
| 19 | 58349.82 | HSS | 90401309031 | Ni | 1200120240 | |
| 20 | 58389.35 | Int | 90401309033 | Ni | 1200120267 | |
| 21 | 58397.30 | LHS | 90401309035 | - | ||
| 22 | 58404.95 | LHS | 90401309037 | - | ||
| 23 | 58420.07 | LHS | 90401309039 | Ni | 1200120290 | |
| 0.99 | ||||||
| MAXI J1535–571 Group 1 | ||||||
| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
| 1 | 58004.35 | LHS | 90401309029 | S | 00010264003 | |
| 2 | 58008.72 | HIMS | 80302309002 | Ni | 1050360104 | |
| 3 | 58010.26 | HIMS | 80402302002 | Ni | 1050360106 | |
| 4 | 58010.57 | HIMS | 80402302004 | - | ||
| 5 | 58013.22 | HIMS | 80302309004 | Ni | 1050360109 | |
| 6 | 58013.48 | HIMS | 80302309006 | Ni | 1050360109 | |
| 7 | 58016.87 | SIMS | 80302309010 | - | ||
| 8 | 58017.21 | SIMS | 80302309010 | - | ||
| : | 1.07 | |||||
| MAXI J1535–571 Group 2 | ||||||
| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
| 9 | 58023.11 | SIMS | 80402302006 | - | ||
| 10 | 58032.71 | SIMS | 80302309012 | - | ||
| 11 | 58034.26 | Int | 80402302008 | - | ||
| 12 | 58048.99 | LHS | 80302309014 | S | 00088245004 | |
| 13 | 58050.98 | LHS | 80402302009 | S | 00088246001 | |
| 14 | 58051.38 | LHS | 80402302010 | - | ||
| 15 | 58090.43 | HSS | 90301327002 | - | ||
| 16 | 58090.77 | HSS | 90301327004 | - | ||
| : | 1.15 | |||||
=-1in
| MAXI J1348–630 | ||||||||||
| Model | Parameter | Epoch 1 | Epoch 2 | Epoch 3 | Epoch 4 | Epoch 5 | Epoch 6 | Epoch 7 | Epoch 8 | Epoch 9 |
| tbabs | nH ( atoms cm-2)555nH, a, i, and AFe tied across all epochs in each group | |||||||||
| diskbb | Tin (keV) | |||||||||
| Nd | ||||||||||
| relxilllp | h (rg) 666-p means lower bound is lower limit,+p means upper bound at upper limit,p means parameter pegged at limit | |||||||||
| a | ||||||||||
| i (degrees) | ||||||||||
| Rin (rg) | ||||||||||
| log | ||||||||||
| AFe | ||||||||||
| Ecut (keV) | ||||||||||
| Nr | ||||||||||
| Constant | FNuB 777Multiplicative factor for NuSTAR FPMA | |||||||||
| FSup888Multiplicative factor for supplemental data, if used. | ||||||||||
| = 1.07 | ||||||||||
=-1in
| MAXI J1820+070 Group 1 | |||||||||||
| Model | Par | Epoch 1 | Epoch 2 | Epoch 3 | Epoch 4 | Epoch 5 | Epoch 6 | Epoch 7 | Epoch 8 | ||
| tbabs | nH ( atoms cm-2) | ||||||||||
| diskbb | Tin (keV) | ||||||||||
| Nd | |||||||||||
| bbody | kT (keV) | ||||||||||
| Nbb | |||||||||||
| relxilllp | h (rg) | ||||||||||
| a | |||||||||||
| i (degrees) | |||||||||||
| Rin (rg) | |||||||||||
| log | |||||||||||
| AFe | |||||||||||
| Ecut (keV) | |||||||||||
| Nr | |||||||||||
| Constant | FNuB | ||||||||||
| FSup | |||||||||||
| = 1.01 | |||||||||||
=-1in
| MAXI J1820+070 Group 1 with Hydrogen Column Density Fixed | |||||||||||
| Model | Par | Epoch 1 | Epoch 2 | Epoch 3 | Epoch 4 | Epoch 5 | Epoch 6 | Epoch 7 | Epoch 8 | ||
| diskbb | nH ( atoms cm-2) | ||||||||||
| Tin (keV) | |||||||||||
| Nd | |||||||||||
| bbody | kT (keV) | ||||||||||
| Nbb | |||||||||||
| relxilllp | h (rg) | ||||||||||
| a | |||||||||||
| i (degrees) | |||||||||||
| Rin (rg) | |||||||||||
| log | |||||||||||
| AFe | |||||||||||
| Ecut (keV) | |||||||||||
| Nr | |||||||||||
| Constant | FNuB | ||||||||||
| FSup | |||||||||||
| = 1.06 | |||||||||||
=-1.0in
| MAXI J1820+070 Group 2 | ||||||||||
| Model | Par | Epoch 9 | Epoch 10 555Disk not detected | Epoch 11 666Disk and blackbody component not detected | Epoch 12 | Epoch 13 | Epoch 14 | Epoch 15 | Epoch 16 | Epoch 17 |
| tbabs | nH ( atoms cm-2) | |||||||||
| diskbb | Tin (keV) | |||||||||
| Nd () | ||||||||||
| bbody | kT (keV) | |||||||||
| Nbb | ||||||||||
| relxilllp | h (rg) | |||||||||
| a | ||||||||||
| i (degrees) | ||||||||||
| Rin (rg) | ||||||||||
| log | ||||||||||
| AFe | ||||||||||
| Ecut (keV) | ||||||||||
| Nr | ||||||||||
| Constant | FNuB | |||||||||
| FSup | ||||||||||
| = 0.98 | ||||||||||
| MAXI J1820+070 Group 3 | |||||||
| Model | Par | Epoch 18 | Epoch 19 | Epoch 20 | Epoch 21 | Epoch 22 | Epoch 23 |
| tbabs | nH ( atoms cm-2) | ||||||
| diskbb | Tin (keV) | ||||||
| Nd | |||||||
| bbody | kT (keV) | ||||||
| Nbb | |||||||
| relxilllp | h (rg) | ||||||
| a | |||||||
| i (degrees) | |||||||
| Rin (rg) | |||||||
| log | |||||||
| AFe | |||||||
| Ecut (keV) | |||||||
| Nr | |||||||
| Constant | FNuB | ||||||
| FSup | |||||||
| = 0.99 | |||||||
| MAXI J1535–571 Group 1 | |||||||||
| Model | Par | Epoch 1 | Epoch 2 | Epoch 3 | Epoch 4 | Epoch 5 | Epoch 6 | Epoch 7 | Epoch 8 |
| tbabs | nH ( atoms cm-2) | ||||||||
| diskbbb | Tin (keV) | ||||||||
| Nd | |||||||||
| relxilllp | h (rg) | ||||||||
| a | |||||||||
| i (degrees) | |||||||||
| Rin (rg) | |||||||||
| log | |||||||||
| AFe | |||||||||
| Ecut (keV) | |||||||||
| Nr | |||||||||
| Constant | FNuB | ||||||||
| FSup | |||||||||
| = 1.07 | |||||||||
| MAXI J1535–571 Group 2 | |||||||||
| Model | Par | Epoch 9 | Epoch 10 | Epoch 11 | Epoch 12 | Epoch 13 | Epoch 14 | Epoch 15 | Epoch 16 |
| tbabs | nH ( atoms cm-2) | ||||||||
| diskbb | Tin (keV) | ||||||||
| Nd | |||||||||
| relxilllp | h (rg) | ||||||||
| a | |||||||||
| i (degrees) | |||||||||
| Rin (rg) | |||||||||
| log | |||||||||
| AFe | |||||||||
| Ecut (keV) | |||||||||
| Nr | |||||||||
| Constant | FNuB | ||||||||
| FSup | |||||||||
| = 1.15 | |||||||||